OpenStudy (anonymous):

the two sides of the triangle are 34 cm and 43 cm , if the included angle is 50 degree. Find the perimeter of the triangle?

OpenStudy (raden):



OpenStudy (raden):

let the third side is x, use the cosine rule to get x. x^2 = 34^2 + 43^2 - 2(34)(43)cos50 solve for x, first

OpenStudy (anonymous):

@RadEn then next step

OpenStudy (raden):

after you get x, then the perimeter is sum all sides perimeter = 34+43+x what's x do you get ?

OpenStudy (anonymous):

x^2 or x?

OpenStudy (raden):

after you get x^2, take the squareroot to get x

OpenStudy (anonymous):

33.55

OpenStudy (anonymous):

correct?

OpenStudy (raden):

i got different like yours

OpenStudy (raden):

x^2 = 34^2 + 43^2 - 2(34)(43)cos50 x^2 = 1156 + 1849 - 2924(cos50) x^2 = 3005 - 2924 * (0.64) x^2 = 3005 - 1871.36 x^2 = 1133.64 x = 33.55 yeah, you are right, LOL

OpenStudy (anonymous):

@RadEn find the length of the smallest altitude of the triangle

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

help please @RadEn

OpenStudy (raden):

still in the figure no. 1 ?

OpenStudy (anonymous):

\[34+43+\sqrt{1156 +1849 - 2924 \text{ Cos}[50{}^{\circ}]}=110.548 \]